OpenStudy (anonymous):

I have to solve for x and y. Please help :) 2x-5y=10 -3x+4y=-15

OpenStudy (anonymous):

multiply the first equation by 3 all the way across multiply the second equation by 2 all the way across then add them the \(x\) terms will add up to zero (cancel, drop out, whatever) and you will be left with only \(y\) to solve for

OpenStudy (anonymous):

\[2x-5y=10\\ -3x+4y=-15 \] \[6x-15y=30\\ -6x+8y=-30 \]

OpenStudy (anonymous):

add and get \(-7y=0\) so \(y=0\)

OpenStudy (anonymous):

THANK YOU <3

OpenStudy (anonymous):

yw don't forget to solve for \(x\)

OpenStudy (anonymous):

Oh yea..

OpenStudy (anonymous):

well shouldnt i solve for that first

OpenStudy (jackellyn):

You could solve for either one first, you just have to go back to solve for the other variable.

OpenStudy (anonymous):

But when I solve for the other variable do I use the other equation too?

OpenStudy (jackellyn):

Yeah, just plug in the answer you got for the variable you already found and then solve for the other variable

OpenStudy (anonymous):

oke

OpenStudy (anonymous):

you can do it with your eyeballs if \(2x-5y=10\) and \(y=0\) you get \(2x=10\) and so \(x=5\)

OpenStudy (anonymous):

:) thanks i completely understand